Dilute sulphuric acid is to be pumped between tow storage tanks using 800m long pipe with 5cm inner diameter at a rate of 100m/hr. the overall difference between the liquid levels in the two tanks is 5cm . the tanks are open to atmospheric through vents . the specific gravity of the acid is 1.3 and the viscous is 0.001Ns/msquare . in the pipe work, there is a globe valve(fully open equivalent to 100pipe diameters ) and four 90 degree standard radius bends (30 pipe diameters each ) the relative roughness of the pipe is 0.002. Any other losses can be considered negligible. what is the power requirement of the centrifugal pump used if the efficiency of the pump is 64% ? take the friction factor of the pipe as 0.031 


1
Expert's answer
2021-04-12T01:46:38-0400

solution.

P= Q(-"\Delta" p)/"\eta" and -"\Delta" p = 4f(l/d)("\rho" u2)/2 and u =Q/A


u = 0.0277/("\pi"/4 x 0.052) = 14,11m/s

-"\Delta" p = 4 x 0.008 (800/0.05)[1840 x 14.112]/2 = 9.378x106


P = [0.0277 x 9.378x106]/0.6 = 4.329 x 105kW


Answer

4.329 x 105kW